# Unveiling the Ignorance of MLLMs: Seeing Clearly, Answering Incorrectly

### v2

@@ C.5 Discussion 2: What Constitutes an Effective Pair of Positive and Negative Samples for Training?
1. Ask if a certain object exists, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the object mentioned in the ques-pos and ans-pos with the visually similar nonexistent object. Ask if the model identifies the replaced object, the question is defined as ques-neg. Rhetorical questions are preferred. Provide the correct answer to ques-neg as ans-neg. 2. Ask a question about the name/category/color/texture/shape/pose of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the name/category/color/texture/shape/pose of the object mentioned in the ques-pos. Ask a question about if the modified name/category/color/texture/shape/pose of an object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 3. Ask a question about the number/position of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the number/position of the object mentioned in the ques-pos with a similar number/position as interference. Ask a question about if the modified number/position of the object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 4. Ask a question about the topic of abstract-knowledge/concrete-knowledge/professional-knowledge in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the knowledge related to the topic mentioned in the ques-pos with similar knowledge as interference. Ask a distractor question about if the modified knowledge related to the topic is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. - Candidate topics of abstract knowledge: emotions, aesthetics, connotations, symbols, culture, news, allusions, legends, common sense, functions, and so on. - Candidate topics of concrete knowledge: landmarks, celebrities, well-known objects, and so on. - Candidate topics of professional knowledge: specific knowledge in various vertical fields, industries, disciplines, and so on. 5. Ask a question about the activity of an object and/or interaction/relationship between objects or persons in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the activity/interaction/relationship mentioned in the ques-pos with a similar activity/interaction/relationship as interference. Ask a distractor question about if the modified activity/interaction/relationship is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 6. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. 7. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as ”in the conversation,” ”mentioned in the conversation,” or ”in the image.” 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely, and not speculative. 5. Do not generate questions that the answer is ”uncertain”. 6. The format of the generated questions and answers is: ’id’: ’json[’id’]’, ’image’: ’json[’image’]’, ’conversations-pos’: [’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’,…, ’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’], ’conversations-neg’: [’from’: ’human’, ’value’: ’ques-neg ’, ’from’: ’gpt’, ’value’: ’ans-neg’,…, ’from’: ’human’, ’value’: ’ques-neg’, ’from’: ’gpt’, ’value’: ’ans-neg’] STEP-3: Output JSON sample
@@
Table 10: Ablation about different combinations of data. ”Version 2/3-Pos.” denotes positive samples generated by the respective prompt version, while ”Version 2/3-Neg.” represents the corresponding negative samples.
@@
Table 11: Ablation about different combinations of data on MMVU test set. ”Version 2/3-Pos.” denotes positive samples generated by the respective prompt version, while ”Version 2/3-Neg.” represents the corresponding negative samples.
@@
Appendix D Training an MLLM on the Proposed MMVU-Train
D.1 Framework.
@@ D.5 Expriment Results
Appendix E Examples of Paired Positive and Negative Samples Generated
@@
Positive Conversation Question: Are there two people in the kitchen? Option: {”Yes, there are two people.”, ”Yes, there are three people.”, ”No, there is only one person.”, ”No, there are four people.”} Answer: Yes, there are two people. Question: Are the people in the kitchen cooking? Option: {”Yes, they are cooking.”, ”No, they are cleaning.”, ”No, they are eating.”, ”Yes, they are fixing something.”} Answer: Yes, they are cooking. Question: How many bowls are there on the counter? Option: {”One bowl.”, ”Four bowls.”, ”Six bowls.”, ”Two bowls.”} Answer: Four bowls. Question: What is the main activity happening? Option: {”Preparing a meal.”, ”Washing dishes.”, ”Playing games.”, ”Reading books.”} Answer: Preparing a meal. Question: Is there a dishwasher in the image? Option: {”Yes, the dishwasher is there.”, ”No, there’s no dishwasher.”, ”It’s unclear.”, ”There is a microwave instead.”} Answer: Yes, the dishwasher is there. Question: Is there a pot on the table? Option: {”Yes, there is a pot.”, ”No, there isn’t a pot.”, ”Yes, there are two pots.”, ”No, there are only pans.”} Answer: Yes, there is a pot. Question: Are the people preparing food in a modern kitchen? Option: {”Yes, they are.”, ”No, it’s an old kitchen.”, ”Yes, in a living room.”, ”No, it’s a commercial kitchen.”} Answer: Yes, they are.
@@
Negative Conversation Question: Are there three people in the kitchen? Option: {”Yes, there are two people.”, ”Yes, there are three people.”, ”No, only one person.”, ”No, four people.”} Answer: No, only one person. Question: Are the people in the kitchen cleaning? Option: {”Yes, they are cooking.”, ”No, they are cleaning.”, ”No, they are eating.”, ”Yes, fixing something.”} Answer: Yes, they are cooking. Question: How many bowls are there on the countertop? Option: {”One bowl.”, ”Four bowls.”, ”Six bowls.”, ”Two bowls.”} Answer: Four bowls. Question: What is the secondary activity happening? Option: {”Preparing a meal.”, ”Washing dishes.”, ”Playing games.”, ”Reading books.”} Answer: Preparing a meal. Question: Is there a microwave in the image? Option: {”Yes, the microwave is there.”, ”No, there’s no microwave.”, ”It’s unclear.”, ”There is a dishwasher.”} Answer: There is a dishwasher. Question: Is there a pan on the table? Option: {”Yes, there is a pot.”, ”No, there isn’t a pot.”, ”Yes, there are two pans.”, ”No, only bowls.”} Answer: Yes, there is a pot. Question: Are the people preparing food in a commercial kitchen? Option: {”Yes, they are.”, ”No, it’s old kitchen.”, ”Yes, a living room.”, ”No, modern kitchen.”} Answer: No, modern kitchen.
@@
Appendix F The Details of “Instruction prompt”.
@@
Appendix G Details about Content Guided Refinement Strategy
G.1 Prompt for extraction visual information
@@
G.2 Examples of Content Guided Refinement Strategy Results
@@ D. Closestool
Appendix H Details about Visual Attention Refinement Strategy
H.1 Detailed Description
@@
H.2 Examples of visual attention refinement strategy
@@
Appendix I Broader Impact and Limitation

### v3

@@ 8.5 Discussion 2: What Constitutes an Effective Pair of Positive and Negative Samples for Training?
1. Ask if a certain object exists, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the object mentioned in the ques-pos and ans-pos with the visually similar nonexistent object. Ask if the model identifies the replaced object, the question is defined as ques-neg. Rhetorical questions are preferred. Provide the correct answer to ques-neg as ans-neg. 2. Ask a question about the name/category/color/texture/shape/pose of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the name/category/color/texture/shape/pose of the object mentioned in the ques-pos. Ask a question about if the modified name/category/color/texture/shape/pose of an object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 3. Ask a question about the number/position of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the number/position of the object mentioned in the ques-pos with a similar number/position as interference. Ask a question about if the modified number/position of the object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 4. Ask a question about the topic of abstract-knowledge/concrete-knowledge/professional-knowledge in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the knowledge related to the topic mentioned in the ques-pos with similar knowledge as interference. Ask a distractor question about if the modified knowledge related to the topic is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. - Candidate topics of abstract knowledge: emotions, aesthetics, connotations, symbols, culture, news, allusions, legends, common sense, functions, and so on. - Candidate topics of concrete knowledge: landmarks, celebrities, well-known objects, and so on. - Candidate topics of professional knowledge: specific knowledge in various vertical fields, industries, disciplines, and so on. 5. Ask a question about the activity of an object and/or interaction/relationship between objects or persons in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the activity/interaction/relationship mentioned in the ques-pos with a similar activity/interaction/relationship as interference. Ask a distractor question about if the modified activity/interaction/relationship is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 6. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. 7. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely, and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’id’: ’json[’id’]’, ’image’: ’json[’image’]’, ’conversations-pos’: [’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’,…, ’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’], ’conversations-neg’: [’from’: ’human’, ’value’: ’ques-neg ’, ’from’: ’gpt’, ’value’: ’ans-neg’,…, ’from’: ’human’, ’value’: ’ques-neg’, ’from’: ’gpt’, ’value’: ’ans-neg’] STEP-3: Output JSON sample
@@
Table 14: Ablation about different combinations of data. "Version 2/3-Pos." denotes positive samples generated by the respective prompt version, while "Version 2/3-Neg." represents the corresponding negative samples.
@@
Table 15: Ablation about different combinations of data on MMVU test set. "Version 2/3-Pos." denotes positive samples generated by the respective prompt version, while "Version 2/3-Neg." represents the corresponding negative samples.
@@
9 Training an MLLM on the Proposed MMVU-Train
9.1 Framework.
@@ 9.5 Expriment Results
10 Examples of Paired Positive and Negative Samples Generated
@@
Positive Conversation Question: Are there two people in the kitchen? Option: {"Yes, there are two people.", "Yes, there are three people.", "No, there is only one person.", "No, there are four people."} Answer: Yes, there are two people. Question: Are the people in the kitchen cooking? Option: {"Yes, they are cooking.", "No, they are cleaning.", "No, they are eating.", "Yes, they are fixing something."} Answer: Yes, they are cooking. Question: How many bowls are there on the counter? Option: {"One bowl.", "Four bowls.", "Six bowls.", "Two bowls."} Answer: Four bowls. Question: What is the main activity happening? Option: {"Preparing a meal.", "Washing dishes.", "Playing games.", "Reading books."} Answer: Preparing a meal. Question: Is there a dishwasher in the image? Option: {"Yes, the dishwasher is there.", "No, there’s no dishwasher.", "It’s unclear.", "There is a microwave instead."} Answer: Yes, the dishwasher is there. Question: Is there a pot on the table? Option: {"Yes, there is a pot.", "No, there isn’t a pot.", "Yes, there are two pots.", "No, there are only pans."} Answer: Yes, there is a pot. Question: Are the people preparing food in a modern kitchen? Option: {"Yes, they are.", "No, it’s an old kitchen.", "Yes, in a living room.", "No, it’s a commercial kitchen."} Answer: Yes, they are.
@@
Negative Conversation Question: Are there three people in the kitchen? Option: {"Yes, there are two people.", "Yes, there are three people.", "No, only one person.", "No, four people."} Answer: No, only one person. Question: Are the people in the kitchen cleaning? Option: {"Yes, they are cooking.", "No, they are cleaning.", "No, they are eating.", "Yes, fixing something."} Answer: Yes, they are cooking. Question: How many bowls are there on the countertop? Option: {"One bowl.", "Four bowls.", "Six bowls.", "Two bowls."} Answer: Four bowls. Question: What is the secondary activity happening? Option: {"Preparing a meal.", "Washing dishes.", "Playing games.", "Reading books."} Answer: Preparing a meal. Question: Is there a microwave in the image? Option: {"Yes, the microwave is there.", "No, there’s no microwave.", "It’s unclear.", "There is a dishwasher."} Answer: There is a dishwasher. Question: Is there a pan on the table? Option: {"Yes, there is a pot.", "No, there isn’t a pot.", "Yes, there are two pans.", "No, only bowls."} Answer: Yes, there is a pot. Question: Are the people preparing food in a commercial kitchen? Option: {"Yes, they are.", "No, it’s old kitchen.", "Yes, a living room.", "No, modern kitchen."} Answer: No, modern kitchen.
@@
11 The Details of “Instruction prompt”.
@@
12 Details about Content Guided Refinement Strategy
12.1 Prompt for extraction visual information
@@
12.2 Examples of Content Guided Refinement Strategy Results
@@ D. Closestool
13 Details about Visual Attention Refinement Strategy
13.1 Detailed Description
@@
13.2 Examples of visual attention refinement strategy
@@
14 Broader Impact and Limitation